### v1

# On the Lie algebra of a Malcev algebra

Abstract. This paper develops the structure theory of a Malcev algebra via the consideration of its most important and largest Lie (sub-) algebra. We introduce the notion of a Lie algebra which uniquely corresponds to a Malcev algebra and use this correspondence to derive some basic properties of some types of ideals in the Malcev algebra. We then prove the exact nature of the root-space decomposition of a Malcev algebra.
[…]
Keywords: Nonassociative algebra; $J-$Nucleus: Malcev algebra: Lie algebra
§$\bf{\S 1.}\;$ Introduction.
One of the most understood and well-developed class of algebras in the whole of mathematics is the class of Lie algebras. A Lie algebra is generally defined as an algebra $\mathfrak{L}$ (over $\mathbb{R}$ or $\mathbb{C}$) whose product, [,],[$\cdot$,$\cdot$],[ ⋅ , ⋅ ] , is anti-commutative (i.e., $[x,x]=0$) and in which the Jacobian,
$[[x,y],z]+[[y,z],x]+[[z,x],y]=:J(x,y,z),$
vanishes, for all $x,y,z\in\mathfrak{L}.$ Numerous articles, monographs and textbooks by notable authors have been published detailing different aspects of a Lie algebra, their inter-relationships and applications to the life sciences. We shall refer to Varadarajan [10.][10.][ 10 . ] for orientation. Perhaps, the most singular reason for the extreme utility of a Lie algebra, over and above other algebras, stems out of the fact that every Lie algebra corresponds to a Lie group. This affords results to be exchanged between a Lie algebra (with its infinitesimal theory) and its corresponding Lie group (with its topological and manifold structures), with one enriching the other in a uniquely determined way. Indeed no other class of algebras has had the privilege of the class of Lie algebras.
Since the Malcev algebra (also called the Moufang-Malcev algebra) was introduced (See Malcev [5.],[5.],[ 5 . ] , Schafer [9.][9.][ 9 . ] and Paal [6.][6.][ 6 . ]) as a tangent space of an analytic Moufang loop and since it was shown that every Lie algebra is a Malcev algebra, mathematicians have been seeking the Lie-algebra type of approach to the understanding of a Malcev algebra. It is believed that a Malcev algebra is a higher version of a Lie algebra in which $J=J(x,y,z)$ is not necessarily zero. Aspects of the structure theory of a Malcev algebra had been discussed by A. A. Sagle in his University of California thesis [8.].[8.].[ 8 . ] . It had also been found that it is possible to have a form of the Poincare-Birkhoff-Witt Theorem (known for a Lie algebra) for a Malcev algebra, as well as its (nonassociative) universal enveloping algebra (See [2.],[2.],[ 2 . ] , [7.][7.][ 7 . ] and [9.][9.][ 9 . ]). All these partial results ([3.][3.][ 3 . ]) were however not enough to give the class of Malcev algebras the esteemed status occupied by one of its sub-classes, the (sub-) class of Lie algebras.
The present paper gives a novel approach to the study of a Malcev algebra via the consideration of a distinguished Lie (sub-) algeba closely attached to it. The paper is organised as follows: §2 considers the structure theory of a Malcev algebra while §develops its basic ideal theory. The root-space decomposition of a Malcev algebra is the subject of §4.
§$\bf{\S 2.}$ Structure theory of a Malcev algebra
We shall start by giving a precise definition to a Malcev algebra and then give a rapid presentation of aspects of its structure.
2.1 Definition. A Malcev algebra $A$ over a field $k$ is any nonassociative algebra over $k$ endowed with a distributive multiplication [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] satisfying the relations
$[a,a]=0$
and
$[J(a,b,c),a]=J(a,b,[a,c]),$
for all $a,b,c\in A.\;\Box$
It should be clear from the above definition that a Lie algebra is in particular a Malcev algebra: Indeed, $[J(a,b,c),a]=[0,a]=0=J(a,b,[a,c]).$ One could then say that a Malcev algebra is a Lie algebra whose Jacobian may not necessarily vanish but satisfies $[J(a,b,c),a]=J(a,b,[a,c]).$ This relationship assures us of many examples of a Malcev algebra while examples of a Malcev non-Lie algebra could be found in the literature, Sagle [8.].[8.].[ 8 . ] .
The result that got the subject started is the following.
2.2 Lemma. ([8.],[8.],[ 8 . ] , p. $435$) Let $A$ denote a Malcev algebra and let $J(A,A,A)$ denote its linear subspace spanned by all elements of the form $J(a,b,c)$ (for $a,b,c\in A$). Then $J(A,A,A)$ is an ideal of $A\;\Box.$
We shall often simply write $J$ for the ideal $J(A,A,A).$ It should be noted that for a Lie algebra, $A,$ the ideal $J(A,A,A)$ is $\{0\}.$ It then follows that the residue-class algebra $A/J$ is well defined. In the special case of a Lie algebra $A$ we have that
$A/J=A/\{0\}\simeq A.$
It happens that $J(A,A,A)$ is the most important ideal for a Malcev non-Lie algebra. Indeed, we have the following.
2.3 Lemma. ([8.],[8.],[ 8 . ] , p. $435$) Let $A$ denote a Malcev algebra of characteristic not $2$ or $3$ Then $J(A,A,A)$ is the smallest ideal of $A$ for which $A/J$ is a Lie algebra..\;$\Box$. □
This in particular shows that $J$ is uniquely determined as an ideal of $A.$ Another distinguished subset of a Malcev algebra is defined in the following.
2.4 Definition. Let $A$ denote a Malcev algebra. The $J-$nucleus of $A$ is given as
$N=\{x\in A:J(x,A,A)=0\}.\;\Box$
This means that $N$ is the maximal subset of $A$ for which $J(N,A,A)=0$ and may also be shown to an ideal of $A.$ Hence $N$ is uniquely determined as an ideal of $A.$ These two uniqely determined ideals $J$ and $N$ of the Malcev algebra $A$ annihilates each other.
2.5 Lemma. ([8.],[8.],[ 8 . ] , p. $440$) If $A$ is a Malcev algebra, then $NJ=0.\;\Box$
The third most important ideal of a Malcev algebra is also given below.
2.6 Definition. Let $A$ denote a Malcev algebra and define $\triangle(x,y)$ as the element of $A$ for which
$[z,\triangle(x,y)]=J(z,x,y),$
for all $x,y,z\in A.$
Elementary properties of $\triangle$ may be found in Sagle[8.][8.][ 8 . ] while the crucial result concerning $\triangle$ is the following.
2.7 Theorem. ([8.],[8.],[ 8 . ] , p. $432$) Let $A$ denote a Malcev algebra of characteristic not $2$ or $3$ and let $\triangle(A,A)$ denote the linear subspace of $A$ spanned by all the $\triangle(x,y)$ (for all $x,y\in A$). Then $\triangle(A,A)$ is a Lie algebra under commutation..\;$\Box$. □
The last result shows that $\triangle(A,A)$ is not a Lie sub-algebra of $A,$ since it is only a Lie algebra with respect to the commutation
$(x,y)=[x,y]-[y,x],$
$x,y\in A.$ Hence, it is bound to exhibit properties not known to $A.$ Moreover, by Proposition $5.8$ of Sagle [8.],[8.],[ 8 . ] , the Lie algebra, $\triangle(A,A)$ is not uniquely determined from $A.$
Our approach in this paper is to employ the well-known fine structure of the Lie (sub-)algebra $N$ of a Malcev algebra $A$ to get across to all of $A$ via the residue-class Lie algebra $A/J.$ The first major result along this line is the following.
2.8 Theorem. ([8.],[8.],[ 8 . ] , p. $442$) Let $A$ denote a Malcev algebra such that $\triangle(A,A)$ is completely reducible in $A,$ then
$A=N\bigoplus J,$
where $N$ is completely reduced by $\triangle(A,A)$ and $J$ is a semi-simple ideal of $A.\;\Box$
The hypotheses of Theorem $1.8$ hold if $A$ is semi-simple. This theorem gives a very useful relationship between the uniquely determined ideals $N$ and $J$ (of $A$) in the light of the residue-class Lie algebra $A/J.$ This is our point of departure in this paper.
§$\bf{\S 3.}$ Ideal theory of a Malcev algebra
We use the truth of Theorem $2.8$ to define the Lie algebra of a Malcev algebra.
3.1 Definition. Let $A$ denote a Malcev algebra. The Lie algebra corresponding to or of $A$ is defined as the residue-class (Lie algebra) $A/J.\;\Box$
3.2 Lemma. Every Malcev algebra has a unique Lie algebra corresponding to it.
Proof. Let $A$ denote any Malcev algebra, then the $J-$nucleus $N$ is non-empty and it is the maximal subset of $A$ for which $J(N,A,A)=0.$ The direct sum $A=N\bigoplus J$ of Theorem $2.8$ thus implies that $A/J\simeq N.$ These respectively imply the existence and uniqueness of the Lie algebra $A/J.\;\Box$
In the case where $A$ is a semi-simple Malcev algebra, its corresponding Lie algebra $A/J$ is equally semi-simple. We shall denote by $\phi$ the canonical epimorphism, $A:\rightarrow A/J$ and shall often make use of the following basic result of commutative algebra.
3.3 Lemma. ([1.],[1.],[ 1 . ] , p. $2$) Let $A$ denote any Malcev algebra. There is a one-to-one order preserving correspondence between the ideals $\mathfrak{b}$ of $A$ which contain $J$ and the ideals $\overline{\mathfrak{b}}$ of the Lie algebra $A/J,$ given by $\mathfrak{b}=\phi^{-1}(\overline{\mathfrak{b}}),$ in which prime ideals correspond to prime ideals..\;$\Box$. □
It therefore follows from the last lemma that, even though there are examples to show that it is not always true that the product of ideals of a MAlcev algebra $A$ is an ideal of $A$ (which is in contrast to the fine property of ideals in a Lie algebra), this failure could now be given a proof and could also be salvaged for ideals of $A$ contain $J$ as contained as follows.
3.4 Proposition. If $\mathfrak{a}$ and $\mathfrak{b}$ are any two ideals of a Malcev algebra, $A,$ both of which contain $J,$ then $\mathfrak{a}\mathfrak{b}$ is an ideal of $A.$
Proof. The one-to-one correspondence of the last lemma implies that $\mathfrak{a}=\phi^{-1}(\overline{\mathfrak{a}})$ and $\mathfrak{b}=\phi^{-1}(\overline{\mathfrak{b}}),$ for some ideals $\overline{\mathfrak{a}}$ and $\overline{\mathfrak{b}}$ in the Lie algebra $A/J.$ Thus
$\mathfrak{a}\mathfrak{b}=\phi^{-1}(\overline{\mathfrak{a}})\phi^{-1}(\overline{\mathfrak{b}})=\phi^{-1}(\overline{\mathfrak{a}}\overline{\mathfrak{b}}),$
where it is known that $\overline{\mathfrak{a}}\overline{\mathfrak{b}}$ is an ideal in $A/J\simeq N.$ The crucial Lemma $3.3$ thus implies that $\mathfrak{a}\mathfrak{b}$ is an ideal in $A$ containing $J.\;\Box$
Our approach in this paper will further show that the true correspondence between a Malcev algebra and a Lie algebra is via the process of its residue-class Lie algebra $A/J,$ which will make the beautiful results of a Lie algebra available to the residue-class (Lie) algebra and then to the Malcev algebra. This could be gleaned through the following commutative diagram:
$A\rightarrow^{\phi}A/J\rightarrow^{i}N,$
where $\phi$ is the canonical map of Lemma $3.3$ and $i$ is the isomorphism of Lemma $3.2$ It then means that we have the projection
$(i\circ\phi):A\rightarrow N.$
Clearly, $i(N+J)=N.$
The above result, for ideals of $A$ containing $J,$ may now be extended to include any subset of the Lie algebra $A/J\simeq N,$ which would buttress our point of view that it is most profitable to approach the study of a Malcev algebra via its projection map $(i\circ\phi):A\rightarrow N,$ into the Lie algebra $N.$
3.5 Lemma. Let $\mathfrak{h}$ denote a subalgebra (respectively, an ideal) of $N,$ then $H:=(i\circ\phi)(\mathfrak{h})$ is a subalgebra (respectively, an ideal) of $A.$
Proof. The surjectivity of $\phi$ implies that for any $n\in N$ there is $a\in A$ such that $a=(i\circ\phi)(n).\;\Box$
Now that we have a direct access between $A$ and its Lie algebra, $N,$ it would be instructive if the extra condition (included in Proposition $3.4$) of an ideal containing $J$ can be removed. This now takes us to the consideration of a special class of ideals of $N$ meant for this purpose.
3.6 Definition. Let $A$ denote a Malcev algebra. A subset $X$ of the Lie algebra $N$ is called an $i-$ideal of $N$ whenever
$i(n)X\subseteq X,$
for every $n\in A/J.\;\Box$
Lemma $3.5$ (together with Lemma $3.3$) suggests that the requirement of an $i-$ideal $X$ of $N$ is equivalent to having $(i\circ\phi)(a)X\subseteq X,$ for every $a\in A.$ The following result shows the central importance of $i-$ideal of the Lie algebra $N$ in getting access to and in the understanding of all the ideals of a Malcev algebra.
3.7 Theorem. Let $A$ denote any Malcev algebra. Then $\mathfrak{a}$ is an ideal of $A$ iff $(i\circ\phi)(\mathfrak{a})$ is an $i-$ideal of $N.$
Proof. Let $a\in A/J$ and let $y\in(i\circ\phi)(\mathfrak{a}),$ then $y=(i\circ\phi)(x),$ for some $x\in\mathfrak{a},$ an ideal of $A.$ It is clear (by Lemma $3.5$) that $X:=(i\circ\phi)(\mathfrak{a})$ is an additive abelian subgroup of $N$ and that $(i\circ\phi)(ax)\subseteq X.$ Note also that
$(i\circ\phi)(ax)=i(\phi(ax))=i(a\phi(x))=i(a)(i\circ\phi)(x)=i(a)y\subseteq i(a)X$
and that
$i(a)(i\circ\phi)(x)=(i\circ\phi)(ax)\subseteq X.$
Hence, $i(a)X\subseteq X,$ for every $a\in N.$
Conversely, let $\overline{\mathfrak{a}}$ denote an $i-$ideal of $N.$ That is, let $i(a)\overline{\mathfrak{a}}\subseteq\overline{\mathfrak{a}},$ for every $a\in A/J.$ Consider now a subset $\mathfrak{a}$ of $A$ as
$\mathfrak{a}:=\{(\phi^{-1}\circ i^{-1})(y):y\in\overline{\mathfrak{a}}\}.$
Clearly, $(i\circ\phi)(\mathfrak{a})=\overline{\mathfrak{a}}.$ Now let $r\in A$ and let $x\in\mathfrak{a},$ then $x=(\phi^{-1}\circ i^{-1})(y),$ for some $y\in\overline{\mathfrak{a}}.$ Considering the element $rx,$ we have that $(i\circ\phi)(rx)=i(\phi(rx))=i(r\phi(x))=i(r)(i\circ\phi)(x)\in i(r)(i\circ\phi)(\mathfrak{a})\in i(r)\overline{\mathfrak{a}}\subseteq\overline{\mathfrak{a}},$ for all $r\in A.$ In particular, for all $r\in N.$ Hence, $(i\circ\phi)(rx)=i(\phi(rx))\in\overline{\mathfrak{a}},$ for all $r\in N.$ Whence, $rx\in(\phi^{-1}\circ i^{-1})(\overline{\mathfrak{a}})=\mathfrak{a}.\;\Box$
The last Theorem sets up a one-to-one correspondence between (all of) the ideals of a Malcev algebra $A$ and the $i-$ideals of its Lie algebra $N.$ It equally confirms that the Lie algebra $N$ is replete with different types of ideals, while the type which is directly concerned with a Malcev algebra is the $i-$ideals of $N.$ We now give a general proof (in contrast to the usual counter-examples) of a major difference on the behaviour of ideals in a Malcev algebra and in a Lie algebra. The next two results show that this negative behaviour of ideals in a Malcev algebra is due to the properties of $i-$ideals in $N.$
3.8 Proposition. The product of two $i-$ideals of $N$ may not always be an $i-$ideal of $N.$
Proof. Every $i-$ideal of $N$ is in particular an ideal in $A.$ Hence, if $\overline{\mathfrak{a}}$ and $\overline{\mathfrak{b}}$ are $i-$ideals of $N$ they also satisfy $\overline{\mathfrak{a}}\overline{\mathfrak{b}}\subseteq\overline{\mathfrak{a}}$ and $\overline{\mathfrak{a}}\overline{\mathfrak{b}}\subseteq\overline{\mathfrak{b}}.$ Thus
$i(a)(\overline{\mathfrak{a}}\overline{\mathfrak{b}})\subseteq i(a)\overline{\mathfrak{a}}\subseteq\overline{\mathfrak{a}}$
and
$i(a)(\overline{\mathfrak{a}}\overline{\mathfrak{b}})\subseteq i(a)\overline{\mathfrak{b}}\subseteq\overline{\mathfrak{b}},$
for any $a\in A/J.$ Therefore,
$i(a)(\overline{\mathfrak{a}}\overline{\mathfrak{b}})\subseteq\overline{\mathfrak{a}}\bigcap\overline{\mathfrak{b}}\nsubseteq\overline{\mathfrak{a}}\overline{\mathfrak{b}}.\;\Box$
The following may therefore be seen as the first known general result on the product of ideals of a Malcev algebra $A.$
3.9 Theorem. Product of ideals of a Malcev algebra $A$ is not always an ideal of $A.$
Proof. We simply combine Theorem $3.7$ with Proposition $3.8.\;\Box$
Since there is a one-to-one correspondence between ideals of $A$ with the $i-$ideals of the Lie algebra $N,$ it follows that among the $i-$ideals $\overline{\mathfrak{a}}$ of $N$ are those (that form a special class) that corresponds to the ideals $\phi^{-1}(\overline{\mathfrak{a}})$ of $A$ that contain $J$ and therefore satisfy Proposition $3.4$ Basic properties of operations on ideals suggest that we seek these (special ones) among the co-prime $i-$ideals of $N$ in order to complete the proof of Proposition $3.8$
3.10 Lemma. The product of co-prime $i-$ideals of $N$ is an $i-$ideal of $N.$
Proof. For every $a\in N$ and for co-prime $i-$ideals $\overline{\mathfrak{a}}$ and $\overline{\mathfrak{b}}$ of $N,$
$i(a)(\overline{\mathfrak{a}}\overline{\mathfrak{b}})\subseteq\overline{\mathfrak{a}}\bigcap\overline{\mathfrak{b}}\subseteq\overline{\mathfrak{a}}\overline{\mathfrak{b}}.\;\Box$
It is very interesting to confirm if the co-prime $i-$ideals of $N$ correspond to ideals of $A$ containing $J.$
§4. Root-space decomposition of a Malcev algebra
We now consider how the root-space decomposition of the Lie algebra $N$ enriches $A.$
4.1 Definition. Let $A$ denote a Malcev algebra and let $N$ be its corresponding Lie algebra. The adjoint action of $A$ on $N$ is the map
$ad_{A\times N}:A\times N\rightarrow N$
$:(a,n)\rightarrowtail ad_{A\times N}(a,n)$
$=:ad_{A\times N}(a)(n)$
defined as
$ad_{A\times N}(a)(n)=i(ad_{A/J}\phi(a)(i^{-1}(n))).$
The map $ad_{A/J}:A/J\rightarrow A/J$ is the adjoint map on the Lie algebra $A/J,$ given as $ad_{A/J}(x)(y)=[x,y].$ The following is immediate.
4.2 Lemma. The adjoint action of $A$ on $N$ is the adjoint map on $N.$
Proof.
$ad_{A\times N}(a)(n)=i(ad_{A/J}\phi(a)(i^{-1}(n)))=i([\phi(a),i^{-1}(n)])=[(i\circ\phi)(a),n].\;\Box$
One of the key results of a Lie algebra is the following.
4.3 Lemma. ([4.],[4.],[ 4 . ] , p. $132$) Let $A$ denote a finite-dimensional Malcev algebra over $\mathbb{C}$ and let $\mathfrak{h}$ be a nilpotent Lie subalgebra of the Lie algebra $N,$ then the generalised weight spaces $N_{\alpha}$ of $N$ relative to $ad_{N}\mathfrak{h}$ satisfy
$(a)$ $N=\bigoplus N_{\alpha},$ where $N_{\alpha}$ is the set
$\{X\in N:(adH-\alpha(H)\cdot 1)^{n}X=0,\;\forall\;H\in\mathfrak{h},\mbox{for some integer}\;n=n(H,X)\geq 0\};$
$(b)$ $\mathfrak{h}\subseteq N_{o};$
$(c)$ $[N_{\alpha},N_{\beta}]\subset N_{\alpha+\beta}$ (with the understanding that $N_{\alpha+\beta}$ is 00 if $\alpha+\beta$ is not a generalised weight).
We then use these properties of $N_{\alpha}$ in $N$ to develop an equivalent theory for generalised weight spaces of a finite-dimensional Malcev algebra, $A.$
4.4 Theorem. Let $A$ denote a finite-dimensional Malcev algebra over $\mathbb{C}$ and let $N_{\alpha}$ be the generalised weight spaces $N_{\alpha}$ of $N$ relative to $ad_{N}\mathfrak{h}$ (given in Lemma $4.3$ above). Choose a subset $H$ of $A$ such that $\mathfrak{h}:=(i\circ\phi)(H)$ is a nilpotent Lie subalgebra of the Lie algebra $N.$ Then the generalised weight spaces $A_{\alpha}$ of $A$ relative to $ad_{A\times N}H$ satisfy
$(a)$ $A=\bigoplus A_{\alpha},$ where $A_{\alpha}$ is given as the set
$\{a\in A:(ad_{A\times N}h-\alpha(h)\cdot 1)^{n}a=0,\;\forall\;h\in\mathfrak{h},\mbox{for some integer}\;n=n(h,a)\geq 0\};$
$(b)$ $H\subseteq A_{o};$
$(c)$ $[A_{\alpha},A_{\beta}]\subset N_{\alpha+\beta}$ (with the understanding that $N_{\alpha+\beta}$ is 00 if $\alpha+\beta$ is not a generalised weight).
Proof. We set $A_{\alpha}=(i\circ\phi)(N_{\alpha}).$ Hence,
$[A_{\alpha},A_{\beta}]=[(i\circ\phi)(N_{\alpha}),(i\circ\phi)(N_{\beta})]=i[\phi(N_{\alpha}),\phi(N_{\beta})]$
$=i[N_{\alpha}+J,N_{\beta}+J]=i([N_{\alpha},N_{\beta}]+J)\subseteq i(N_{\alpha+\beta}+J)=N_{\alpha+\beta},$
since $N_{\alpha+\beta}\subseteq N,$ in which $i(N+J)=N.\;\Box$
There is still a lot of the results of Lie algebras (and Lie groups) that may be proven for a Malcev non-Lie algebra.
§5. References.
[1.]
Atiyah, M. F. and Macdonald, I. G., Introduction to commutative algebra, Addison-Wesley Series in Mathematics, Addison-Wesley, Reading. $1969$
[2.]
Bremner, M. R, Hentzel, I. R., Peresi, L. A., Tvalavadze, M. V., Usefi, H., Enveloping algebras of Malcev algebras, Comment. Math. Univ. Carolin. 51, $2$ ($2010$), p. $157-174$
[3]
Harrathi, F., Mabrouk, S., Ncib, O., Silvestrov, S. Malcev Yang-Baxter equation, weighted $\mathfrak{O}-$operators on Malcev algebras and post-Malcev algebras, arXiv:$2206.03629v1$ [math.RA] ($2022$)
[4.]
Knapp, A. W. Lie Groups Beyond an Introduction, Progress in Mathematics, 140 ($2002$)
[5.]
Malcev, A. I., Analytic loops, Matem. Sb. 78, ($1955$), p. $569-578$
[6.]
Paal, E., Note on analytic Moufang loops, Comment. Math. Univ. Carolin. 45, $2$ ($2004$), p. $349-354$
[7.]
Perez-Izquiredo, J. M., Shestakov, I. P., An envelope for Malcev algebras, J. Algebra $\bf 272$ ($2004$), p. $379-393$
[8.]
Sagle, A. A., Malcev algebras, Trans. Amer. Math. Soc. 101 ($1961$), p. $426-458$
[9.]
Schafer, R. D., An introduction to nonassociative algebras, Academic Press, New York, ($1966$).
[10.]
Varadarajan, V. S., Lie groups, Lie algebras and their representations, Springer-Verlag, New York, ($1984$).

### v2

@@
 The third most important ideal of a Malcev algebra is also given below.
-2.6 Definition. Let $A$ denote a Malcev algebra and define $\triangle(x,y)$ as the element of $A$ for which
+2.6 Definition. Let $A$ denote a Malcev algebra and define $\triangle(x,y)$ by the requirement
 $[z,\triangle(x,y)]=J(z,x,y),$
@@
 ] , p.
-$432$) Let $A$ denote a Malcev algebra of characteristic not $2$ or $3$ and let $\triangle(A,A)$ denote the linear subspace of $A$ spanned by all the $\triangle(x,y)$ (for all $x,y\in A$).
+$432$) Let $A$ denote a Malcev algebra of characteristic not $2$ or $3$ and let $\triangle(A,A)$ denote the linear span of all the $\triangle(x,y)$ (for all $x,y\in A$).
 Then $\triangle(A,A)$ is a Lie algebra under commutation..\;$\Box$.
 □
-The last result shows that $\triangle(A,A)$ is not a Lie sub-algebra of $A,$ since it is only a Lie algebra with respect to the commutation
+The last result shows that $\triangle(A,A)$ is not realized as a Lie sub-algebra of $A,$ since it is only a Lie algebra with respect to the commutation
 $(x,y)=[x,y]-[y,x],$